A rectangle has length 3xcm and width 3x/2cm
A) Find the area of rectangle
B) Then find the area if x=12

Answer by oscargut(2103)

A) Area = 3x(3x)/2 =(9x^2)/2 cm^2
B) Area = 9(12)^2/2 = 648 cm^2
I can help you with more questions at: mthman@gmail.com
Thanks